# Color Equivariant Network

### v2

@@ Abstract
Group equivariant convolutional neural networks have been designed for a variety of geometric transformations from 2D and 3D rotation groups, to semi-groups such as scale. Despite the improved interpretability, accuracy and generalizability afforded by these architectures, group equivariant networks have seen limited application in the context of perceptual quantities such as hue and saturation, even though their variation can lead to significant reductions in classification performance. In this paper, we introduce convolutional neural networks equivariant to variations in hue and saturation by design. To achieve this, we leverage the observation that hue and saturation transformations can be identified with the 2D rotation and 1D translation groups respectively. Our hue-, saturation-, and fully color-equivariant networks achieve equivariance to these perceptual transformations without an increase in network parameters. We demonstrate the utility of our networks on synthetic and real world datasets where color and lighting variations are commonplace.
@@ 1 Introduction
Figure 1: Hue equivariance. The equivariance of our hue equivariant model is illustrated by the commutativity of the (hue) rotation and neural network mapping. A hue rotation of 90 in the input image space (top-left to bottom-left), results in a feature map rotation at each layer of the network (top-right to bottom-right).
The tremendous progress of image classification in the last decade can be readily attributed to the development of deep convolutional neural networks [24, 30, 19]. The highly nonlinear mapping and large parameter space does not lend itself to interpretation easily, but even in early networks, representations of geometry and color were observed and recognized for their importance in representation learning [24, 26]. While a relatively large body of literature has worked to improve the robustness of convolutional neural networks to geometric transformations [4, 7, 16, 28, 20, 17], improving their robustness to perceptual variation has garnered considerably less attention.
A commonly used heuristic for improving network robustness to color variation is to perform mean subtraction and normalization on training set examples. This approach can work well when the training and testing datasets are drawn from the same distribution; however, for data that are collected at different points in time, or with different sensors, this not likely to be the case. Consider, for example, the case of medical imaging where images of tissue samples collected from different labs (or from the same lab at different points in time) may have different characteristics due to variability in data collection protocols or imaging processes [33]. This variability presents a significant challenge for convolutional neural networks which have been found to be sensitive to color variations even on the level of individual blocks [11]. Moreover, when presented with color perturbations, conventional networks exhibit a significant drop in classification performance [10]. One approach to mitigate the effect of color variation is to ignore color entirely by enforcing color invariance. This can be done by converting input images to grayscale, or enforcing representation similarity across color as is done in [27]. This approach has its own challenges, however, since in many domains, color is an important cue for classification. Another approach is to use dataset augmentation, a technique that can improve robustness in the presence of known transformations [3]. This approach, however, requires extended training times and is not likely to offer improved interpretability.
In this work, we address the challenge of neural network sensitivity to color variation by leveraging the geometric structure of color in a group equivariant network. Geometric deep learning has gained strong interest in recent years due to its ability to capture information in interpretable and generalizable representations; by using a group equivariant architecture, we inherit these characteristics for the context of color variation. Using this approach, our network retains color information in a structured representation throughout the processing pipeline, thus allowing color information to be used and/or discarded with intentionality.
We represent color in the hue saturation luminance color space, and leverage the insight that the hue, and saturation can be modeled with geometric group structure. Specifically, we identify variations in hue with the 2D rotation group, and variations in saturation with the 1D translation group. With this identification, we propose a group equivariant network where variations in hue and saturation are represented as a geometric transformations.
To summarize, our key contributions are:
  • We introduce networks that are equivariant to shifts in hue and saturation, without requiring an increase in the number of network parameters.
  • Our networks perform on par or better than conventional architectures in the presence of unseen color combinations on synthetic, and real world datasets.
  • Our networks produce interpretable feature representations that are useful for downstream tasks such as color based sorting.
@@ 2 Related work
Group Equivariant Networks
The strong generalization performance of convolutional neural networks on image processing tasks is attributed, in part, to the equivariance of planar convolution to 2D translations. This insight has garnered considerable attention and led to a strong interest in the development of convolutional neural networks equivariant to other symmetry groups [22, 9]. Previous works introduce group equivariant convolutional networks for finite and continuous groups. A framework for finite group convolution was developed in [7], and demonstrated for the 2D rotation, and reflection groups. The authors achieve equivariance to specific symmetry groups by convolving input images and feature maps with the group orbit of a learned filter bank. In [36], the authors design for equivariance to the continuous 2D rotation group by constraining filter representations to have circular harmonic structure.
Group equivariant networks have also been designed for use in the context of scale symmetry. In [12], the authors represent input images in log-polar coordinates where 2D rotation and scale transformations present as 2D translations. To navigate the semi-group structure of scale transformations, the authors in [35] approximate the scale space as finite and use dilated convolutions in a group equivariant architecture. The SREN network proposed in [31] designs for equivariance to the continuous 2D rotation and isotropic scaling group by constraining filter representations to be expressed as the linear combination of a windowed Fourier basis.
Group equivariant networks have also been designed for groups acting in higher dimensions. For 3D rotational symmetry,  [32] and [13] use a spherical harmonics based representation with 3D point clouds and spherical image inputs respectively; and  [2] introduce an equivariant network for transformations of the Euclidean group in 3D. In [14], the authors introduce a framework for general Lie groups whose exponential map is surjective, and  [8] introduces a framework for more general manifolds. By designing for known symmetries in the task, these works provide improved interpretability, training efficiency and generalizability over conventional convolutional networks.
Other works encourage transformation equivariance using a soft penalty term on representation dissimilarity [26, 18]. These methods are separate from the group equivariant network literature, and different from what we propose. Our model leverages the geometric structure of hue and saturation to design a hue-saturation equivariant network bringing the benefits of group equivariant networks to perceptual transformations.
@@
Several prior works attempt to mitigate the effect of color variation in image processing tasks, for example, in [6], the authors represent images in a color space where pixel values are invariant to changes in luminance for luminance invariant image segmentation, and in [27], the authors penalize changes in their latent representations due to color variation for color invariant skin lesion identification. While color invariance resolves the effect of color variation, color has been identified as an important cue in neural network representation learning [11, 10] which is discordant with the goal of color invariance. In contrast to these approaches, our work leverages the geometric structure of hue, and saturation to construct a convolutional neural network equivariant to variations in these quantities by design.
Most similar to ours is the model proposed in [29], a hue equivariant network for improved network performance in the presence of color variation. Our color equivariant networks differ from the work proposed in  [29] in three important ways. First, our networks are fully equivariant, in contrast to the hybrid networks proposed in [29]. Second, we lift the input image instead of the filters of the first layer which circumvents the issue of invalid hue rotations suffered by the network in [29]. Finally, we introduce networks equivariant to hue and saturation, which together provide full color equivariance and is unique to our approach.
@@ 3 Preliminaries
The equivariance of our network to hue shifts is illustrated in Figure 1.
@@ 4 Method
Figure 2: Lifting layer. An input image (left) is “lifted” to the hue-saturation group (right) by shifting its hue and saturation values.
@@
In the HSL color space, hue is represented by angular position, and can therefore be identified with the 2D rotation group. As in group equivariant network [7], we consider a finite group representation. Specifically, we identify elements of the discretized hue group, $H_{N}$, with those of the cyclic group $C_{N}$.
@@
$\varphi_{h}(h_{i},x)=((x_{h}+h_{i})(\text{mod}\;255),x_{s},x_{l}),$ (7)
where $(\cdot)\text{mod}(\cdot)$ denotes the modulus operation and is applied to the pixel value rather than its corresponding degree.
An element of the hue group acts on a function on the discrete hue group by the group action $\phi_{h}:H_{N}\times Y\rightarrow Y$, which “rotates” the function on the group. Concretely, for a function $f$ on the discrete hue group $H_{N}$ defined as the concatenation of functions $f=(f_{1},\dots,f_{N})$, the action of an element $h_{i}$ in the hue group $H_{N}$ is given by
@@
In the HSL color space, saturation can be represented by a real number in the interval $[0,1]$. We introduce two approximations to give group structure to the saturation space. First, we observe that there is a bijection between the real numbers and the open interval $(0,1)$, so we use the structure of the group $(\mathbb{R},+)$. Second, we consider a finite subset of the group, a necessary and commonly used practice in both conventional CNNs (recall that the translation group is infinite) and group-equivariant CNNs such as [35].
@@
$\varphi_{s}(s_{i},x)=(x_{h},\text{min}(x_{s}+s_{i},255),x_{l}).$ (9)
An element of the saturation group acts on a function on the discrete saturation group by the group action $\phi_{s}:S_{N}\times Y\rightarrow Y$, which “translates” the function on the group. Concretely, for a function $f$ on the discrete saturation group $S_{N}$ defined as the concatenation of functions $f=(f_{1},\dots,f_{N})$, the action of an element $s_{i}$ in the saturation group $S_{N}$ is given by
@@
An element of the hue-saturation group acts on a function on the discrete saturation group by the group action $\phi_{hs}:H_{N}\times S_{M}\times Y\rightarrow Y$, which “rotates” and “translates” the function on the group. Concretely, for a function $f$ on the discrete hue-saturation group $H_{N}\times S_{M}$ defined as the concatenation of functions $f=(f_{11},\dots,f_{1M},f_{21},\dots,f_{NM})$, the action of an element $(h_{i},s_{j})$ in the hue-saturation group $H_{N}\times S_{M}$ is given by
@@
The first layer of a group convolutional neural network “lifts” the input image to the group (see Figure 2). We can lift an input image to the product space of the image grid $\mathbb{Z}^{2}$, and discretized hue shifts $H_{N}=\{h_{0},h_{1},\dots,h_{N}\}$, by convolving with hue shifted filters,
@@
Here we denote an element of the product space $g_{x,i}$, where the subscript $x$ references an element of $\mathbb{Z}^{2}$, and $j$ references the element $h_{j}$ in $H_{N}$. The input image $f^{0}$ and lifting filters $\psi_{m,n}$ are functions on $\mathbb{Z}^{2}$.
Lifting to the group can also be achieved by transforming the input image rather than the filters,
@@
We use this alternative in our implementation as it is both simple and does not result in pixels falling outside the RBG cube as in [29]. The lifting layers for the saturation group, and hue-saturation group are constructed analogously.
@@
We can show that our lifting layer is equivariant to hue shifts. We let $f^{1}(g_{x,j})=\varphi_{h}(h_{j},f^{0})(x)$ as in equation 14 and show that a hue shift of $f^{0}$ results in a hue shift of $f^{1}$. In the first step we use the fact that $\varphi_{h}$ is a group action (see Appendix C.1), and in the second step we use commutativity of the hue group:
$\varphi_{h}(h_{j},\varphi_{h}(h_{m},f^{0}))(x)$ $=\varphi(h_{j}h_{m},f^{0})(x)$ (15)
@@
For all layers after the first layer, the feature maps $f^{l}$ are on the product space $\mathbb{Z}^{2}\times H_{N}$. Since hue shifts can only be performed on three dimensional inputs, we leverage the identification of hue shifts with the discrete 2D rotation group, i.e., $H_{N}\cong C_{N}$ and interpret the feature maps $f^{l}$ as functions on the product space $G:=\mathbb{Z}^{2}\times C_{N}$. With this interpretation we perform group convolution on $G$ as follows,

### v3

# Learning Color Equivariant Representations

@@ Abstract
In this paper, we introduce group convolutional neural networks (GCNNs) equivariant to color variation. GCNNs have been designed for a variety of geometric transformations from 2D and 3D rotation groups, to semi-groups such as scale. Despite the improved interpretability, accuracy and generalizability of these architectures, GCNNs have seen limited application in the context of perceptual quantities. Notably, the recent CEConv network uses a GCNN to achieve equivariance to hue transformations by convolving input images with a hue rotated RGB filter. However, this approach leads to invalid RGB values which break equivariance and degrade performance. We resolve these issues with a lifting layer that transforms the input image directly, thereby circumventing the issue of invalid RGB values and improving equivariance error by over three orders of magnitude. Moreover, we extend the notion of color equivariance to include equivariance to saturation shift. Our hue-, saturation-, and color-equivariant networks achieve strong generalization to out-of-distribution perceptual variations and improved sample efficiency over conventional architectures. We demonstrate the utility of our approach on synthetic and real world datasets where we consistently outperform competitive baselines.
@@ 1 Introduction
The tremendous progress of image classification in the last decade can be readily attributed to the development of deep convolutional neural networks [1; 2; 3]. The highly nonlinear mapping and large parameter space does not lend itself to interpretation easily, but even in early networks, representations of geometry and color were observed and recognized for their importance [1; 4]. While a relatively large body of literature has worked to improve the robustness of convolutional neural networks to geometric transformations [5; 6; 7; 8; 9; 10], improving their robustness to perceptual variation has garnered considerably less attention.
A commonly used heuristic for improving network robustness to color variation is to perform mean subtraction and normalization on training set examples. This approach can work well when the training and testing datasets are drawn from the same distribution; however, for data that are collected at different points in time, or with different sensors, this is not likely to be the case. Consider, for example, the case of medical imaging where images of tissue samples collected from different labs (or from the same lab at different points in time) may have different characteristics due to variability in data collection protocols or imaging processes [11]. This variability presents a significant challenge for convolutional neural networks which have been found to be sensitive to color variations even on the level of individual blocks [12]. Moreover, when presented with color perturbations, conventional networks exhibit a significant drop in classification performance [13]. One approach to mitigate the effect of color variation is to ignore color entirely by enforcing color invariance. This can be done by converting input images to grayscale, or enforcing representation similarity across color as is done in [14]. This approach has its own challenges, however, since in many domains, color is an important cue for classification. Another approach is to use dataset augmentation, a technique that can improve robustness in the presence of known transformations [15]. This approach, however, requires extended training times and does not provide improved interpretability.
In this work, we address the challenge of neural network sensitivity to color variation by leveraging the geometric structure of color in a group convolutional neural network. Geometric deep learning has gained strong interest in recent years due to its ability to capture information in interpretable and generalizable representations; by using a group convolutional neural network, we inherit these characteristics for the context of color variation. We represent color in the hue saturation luminance (HSL) color space, and leverage the insight that the hue, and saturation can be modeled with geometric group structure. Using this approach, our network retains color information in a structured representation throughout the processing pipeline, thereby allowing color information to be used and/or discarded intentionally.
Recently, Color Equivariant Convolutional Networks [16], proposed a group convolutional neural network architecture for hue equivariant representation learning. Most notably, CEConv identifies hue transformations with the 2D rotation group. However, the proposed lifting layer introduces invalid RGB values which break equivariance, and degrade performance. To resolve this issue, our lifting layer operates on the input image instead of the network filters, which produces an equivariant descriptor without projection induced artifacts. This seemingly minor change improves equivariance error by more than three orders of magnitude, and stabilizes network performance across discretizations. We also expand the notion of color equivariance to capture variation in saturation. Specifically, we identify variations in saturation with the 1D translation group. With this identification, we propose a group equivariant network where variations in hue and saturation are represented as a geometric transformations.
@@
(a) Hue equivariance
(b) Lifting layer
Figure 1: Color equivariant network. (a) The equivariance of our hue equivariant model is illustrated by the commutativity of the (hue) rotation and neural network mapping. A hue rotation of 90 in the input image space (top-left to bottom-left), results in a feature map rotation at each layer of the network (top-right to bottom-right). Corresponding feature maps are highlighted with a blue border. (b) An input image (left) is “lifted” to the hue-saturation group (right) by shifting its hue and saturation values.
@@ 2 Related Work
Group Convolutional Neural Networks
The strong generalization performance of convolutional neural networks on image processing tasks is attributed, in part, to the equivariance of planar convolution to 2D translations. This insight has garnered considerable attention and led to a strong interest in the development of convolutional neural networks equivariant to other symmetry groups [17; 18]. Previous works introduce group convolutional networks for finite and continuous groups. A framework for finite group convolution was developed in [6], and demonstrated for the 2D rotation, and reflection groups. The authors achieve equivariance to specific symmetry groups by convolving input images and feature maps with the group orbit of a learned filter bank. In [19], the authors design for equivariance to the continuous 2D rotation group by constraining filter representations to have circular harmonic structure.
Group convolutional networks have also been designed for use in the context of scale symmetry. In [20], the authors represent input images in log-polar coordinates where 2D rotation and scale transformations present as 2D translations. To navigate the semi-group structure of scale transformations, the authors in [21] approximate the scale space as finite and use dilated convolutions in a group equivariant architecture. The SREN network proposed in [22] designs for equivariance to the continuous 2D rotation and isotropic scaling group by constraining filter representations to be expressed as the linear combination of a windowed Fourier basis.
Group convolutional networks have also been designed for groups acting in higher dimensions. For 3D rotational symmetry, [23] and [24] use a spherical harmonics based representation with 3D point clouds and spherical image inputs respectively; and [25] introduce an equivariant network for transformations of the Euclidean group in 3D. In [26], the authors introduce a framework for general Lie groups whose exponential map is surjective, and [27] introduces a framework for more general manifolds. By designing for known symmetries in the task, these works provide improved interpretability, training efficiency and generalizability over conventional convolutional networks.
Other group equivariance methods encourage transformation equivariance using a soft penalty term on representation dissimilarity [4; 28]. These methods are separate from the group convolutional neural network literature, and different from what we propose. Our model leverages the geometric structure of hue and saturation to design a hue-saturation equivariant network bringing the benefits of group equivariant networks to perceptual transformations.
@@
Several prior works attempt to mitigate the effect of color variation in image processing tasks, for example, in [29], the authors represent images in a color space where pixel values are invariant to changes in luminance for luminance invariant image segmentation, and in [14], the authors penalize changes in their latent representations due to color variation for color invariant skin lesion identification. While color invariance resolves the effect of color variation, color has been identified as an important cue in neural network representation learning [12; 13] which is discordant with the goal of color invariance. In contrast to these approaches, our work leverages the geometric structure of hue, and saturation to construct a convolutional neural network equivariant to variations in these quantities by design.
Most similar to ours is the model proposed in [16], a hue equivariant network for improved network performance in the presence of color variation. Our color equivariant networks differ from the work proposed in  [16] in two important ways. First, we lift the input image instead of the filters of the first layer which circumvents the issue of invalid hue rotations suffered by the network in [16]; and second, we expand the notion of color equivariance by introducing networks that are equivariant to both hue and saturation.
@@ 3 Preliminaries
The equivariance of our network to hue shifts is illustrated in Figure 1(a).
@@ 4 Method
In the HSL color space, hue is represented by angular position, and can therefore be identified with the 2D rotation group. As in group convolutional network [6], we consider a finite group representation. Specifically, we identify elements of the discretized hue group, $H_{N}$, with those of the cyclic group $C_{N}$.
@@
$\varphi_{h}(h_{i},x)=((x_{h}+h_{i})(\text{mod}\;c),x_{s},x_{l}),$ (7)
where $(\cdot)\text{mod}(\cdot)$ denotes the modulus operation and is applied to the pixel value which ranges from 0 to $c$.
An element of the hue group acts on a function on the discrete hue group by the group action $\phi_{h}:H_{N}\times Y\rightarrow Y$, which “rotates" the function on the group. Concretely, for a function $f$ on the discrete hue group $H_{N}$ defined as the concatenation of functions $f=(f_{1},\dots,f_{N})$, the action of an element $h_{i}$ in the hue group $H_{N}$ is given by
@@
The action of $\varphi_{h}$ on an input image, and $\phi_{h}$ on a feature map are shown in Figure 6(a). We prove $\varphi_{h}$ and $\phi_{h}$ are group actions in Appendix A.
@@
We expand the notion of color equivariance proposed in [16] to include equivariance to saturation shifts. In the HSL color space, saturation can be represented by a real number in the interval $[0,1]$; we introduce two approximations to give the saturation space group structure. First, we observe that there is a bijection between the real numbers and the open interval $(0,1)$, so we can use the structure of the group $(\mathbb{R},+)$. Second, we consider a finite subset of the group, a necessary and commonly used practice in both conventional CNNs (recall that the translation group is infinite) and group-equivariant CNNs such as [21]. With these approximations, we identify elements of the discretized saturation group $S_{N}$, with those of the integers with addition, $(\mathbb{Z},+)$.
@@
$\varphi_{s}(s_{i},x)=(x_{h},\text{min}(x_{s}+s_{i},c),x_{l}).$ (9)
An element of the saturation group acts on a function on the discrete saturation group by the group action $\phi_{s}:S_{N}\times Y\rightarrow Y$, which “translates" the function on the group. Concretely, for a function $f$ on the discrete saturation group $S_{N}$ defined as the concatenation of functions $f=(f_{1},\dots,f_{N})$, the action of an element $s_{i}$ in the saturation group $S_{N}$ is given by
@@
The action of $\varphi_{s}$ on an input image, and $\phi_{s}$ on a feature map are shown in Figure 2. We prove $\varphi_{s}$ and $\phi_{s}$ are group actions in Appendix A.
Figure 2: Saturation equivariant feature maps. We illustrate the equivariance of our saturation-equivariant model. A saturation shift in the input image space (top-left to bottom-left), results in a feature map translation at each layer of the network (top-right to bottom-right). Corresponding feature maps are highlighted with a blue border.
@@
An element of the hue-saturation group acts on a function on the discrete saturation group by the group action $\phi_{hs}:H_{N}\times S_{M}\times Y\rightarrow Y$, which “rotates" and “translates" the function on the group. Concretely, for a function $f$ on the discrete hue-saturation group $H_{N}\times S_{M}$ defined as the concatenation of functions $f=(f_{11},\dots,f_{1M},f_{21},\dots,f_{NM})$, the action of an element $(h_{i},s_{j})$ in the hue-saturation group $H_{N}\times S_{M}$ is given by
@@
The first layer of a group convolutional neural network “lifts" the input image to the group (see Figure 1(b)). We can lift an input image to the product space of the image grid $\mathbb{Z}^{2}$, and discretized hue shifts $H_{N}=\{h_{0},h_{1},\dots,h_{N}\}$, by convolving with hue shifted filters,
@@
Figure 3: Impact of cardinality on hue rotation invertibility. Our lifting layer (blue) operates on HSL input images where each hue rotation is invertible. The lifting layer proposed in CEConv (orange) operates on RGB filters and suffers from invalid hue rotations for all discretizations of the hue group except for $N=1$ and $N=3$ (i.e., symmetries of the axis-aligned RGB cube). (Left) We show the impact of invalid hue rotations on a four pixel image. We rotate the image $60^{\circ}$ using our proposed lifting layer (top-left) and the CEConv lifting layer (bottom-left). Subsequently applying a $-60^{\circ}$ rotation yields an image that is indistinguishable from the original using our approach (top-left), and one with visible artifacts using the CEConv approach (bottom-left). (Right) We show the average restored image error for both approaches. Our approach results in a consistently negligible restored image error, however, the CEConv approach results in a restored image error exceeding $7\%$ for all discretizations of the hue group except $N=1$ and $N=3$.
Here we denote an element of the product space $g_{x,i}$, where the subscript $x$ references an element of $\mathbb{Z}^{2}$, and $j$ references the element $h_{j}$ in $H_{N}$. The input image $f^{0}$ and lifting filters $\psi_{m,n}$ are functions on $\mathbb{Z}^{2}$. This is the approach taken in CEonv [16]. The authors shift the hue of a filter in the RGB space by rotating its values about an axis passing through the point $p=(1,1,1)$ as shown in Figure 8 in Appendix B. As described in [16], this approach results in invalid hue rotations for all discretizations of the hue group that are not symmetries of the axis-aligned RGB cube (i.e., $N=1$ and $N=3$). To remedy this, the authors project invalid RGB values to the nearest point on the RGB cube. However, this projection breaks invertibility of the hue shift (see Figure 3) yielding a descriptor that is only approximately hue equivariant.
Lifting to the group can alternatively be achieved by transforming the input image rather than the filters,
@@
We use this approach in our implementation as it circumvents the issue of invalid hue rotations, producing a reliably equivariant descriptor. The lifting layers for the saturation group, and hue-saturation group are constructed analogously.
@@
We can show that our lifting layer is equivariant to hue shifts. We let $f^{1}(g_{x,j})=\varphi_{h}(h_{j},f^{0})(x)$ as in equation 14 and show that a hue shift of $f^{0}$ results in a hue shift of $f^{1}$. In the first step we use the fact that $\varphi_{h}$ is a group action (see Appendix A.1), and in the second step we use commutativity of the hue group:
$\varphi_{h}(h_{j},\varphi_{h}(h_{m},f^{0}))(x)$ $=\varphi_{h}(h_{j}h_{m},f^{0})(x)$ (15)
@@
For all layers after the first layer, the feature maps $f^{l}$ are on the product space $\mathbb{Z}^{2}\times H_{N}$. Since hue shifts can only be performed on three dimensional inputs, we leverage the identification of hue shifts with the discrete 2D rotation group, i.e., $H_{N}\cong C_{N}$ and interpret the feature maps $f^{l}$ as functions on the product space $G:=\mathbb{Z}^{2}\times C_{N}$. With this interpretation we
Figure 4: Model sample efficiency. We show the error improvement (higher is better) over the Z2CNN baseline as a function of the percentage of training examples used. The advantage of our Hue-$N$ models increases as the percentage of training examples used decreases.
Table 1: Generalization to global hue-shift. Classification error on the Hue-shift MNIST dataset is reported. Our hue-equivariant models (Hue-$N$) perform comparably to the conventional CNN model (Z2CNN) on the in-distribution case ($A/A$), and significantly outperform the conventional CNN model on the out-of-distribution case ($A/B$).
Network $A/A$ $A/B$ Params
Z2CNN 1.54 (0.10) 57.38 (22.06) 22,130
Hue-3* 1.79 (0.25) 1.81 (0.29) 22,658
Hue-4* 1.97 (0.25) 2.08 (0.15) 25,690
perform group convolution on $G$ as follows,